OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -4) with a slope of 1/2.

OpenStudy (da_scienceman):

the general form of the equation of a line y-y1=m(x -x1). here, y1 = -4, x1 = 2 while m = 1/2. Can u now write down the said equation of the line

OpenStudy (anonymous):

Can you explain to me how to do it, cause I don't understand it